# Rotational Taylor dispersion in linear flows

### v1

@@ Abstract
The coupling between advection and diffusion in position space can often lead to enhanced mass transport compared to diffusion without flow. An important framework used to characterize the long-time diffusive transport in position space is the generalized Taylor dispersion theory. In contrast, the dynamics and transport in orientation space remains less developed. In this work, we develop a rotational Taylor dispersion theory that characterizes the long-time orientational transport of a spheroidal particle in linear flows that is constrained to rotate in the velocity-gradient plane. Similar to Taylor dispersion in position space, the orientational distribution of axisymmetric particles in linear flows at long times satisfies an effective advection-diffusion equation in orientation space. Using this framework, we then calculate the long-time average angular velocity and dispersion coefficient for both simple shear and extensional flows. Analytic expressions for the transport coefficients are derived in several asymptotic limits including nearly-spherical particles, weak flow, and strong flow. Our analysis shows that at long times the effective rotational dispersion is enhanced in simple shear and suppressed in extensional flow. The asymptotic solutions agree with full numerical solutions of the derived macrotransport equations and results from Brownian dynamics simulations. Our results show that the interplay between flow-induced rotations and Brownian diffusion can fundamentally change the long-time transport dynamics.
@@
colloids, suspensions
@@ 1 Introduction
Transport and mixing of solutes or particles in the presence of hydrodynamics flows are important for various biological and industrial processes. For micron-sized particles immersed in flows, the coupling between advection and diffusion can often lead to enhanced mass transport as compared to diffusion without flow. A classical example of such a coupling effect is the Taylor dispersion of Brownian solutes in pressure-driven channel flows (Taylor, 1953, 1954a, 1954b; Aris, 1956). Brownian motion allows the solute particles to migrate across streamlines, and then be advected downstream with different velocities. At long times, the coupling between transverse diffusion and longitudinal advection gives rise to diffusive transport of the solutes with an effective longitudinal dispersion coefficient that can be much larger than the bare diffusivity of the solute particle. Since the work of Taylor (1953), a generalized Taylor dispersion (GTD) framework (Frankel & Brenner, 1989) has been developed to accommodate a wide range of transport problems including complex geometries, chemical reactions, spatial and/or time periodicity, and active particles (Brenner, 1980; Shapiro & Brenner, 1986, 1987, 1990; Morris & Brady, 1996; Hill & Bees, 2002; Bearon, 2003; Manela & Frankel, 2003; Zia & Brady, 2010; Takatori & Brady, 2014; Burkholder & Brady, 2017; Alonso-Matilla et al., 2019; Jiang & Chen, 2019; Peng & Brady, 2020).
@@
Using experiments and particle-based simulations, Leahy et al. (2013) showed that advection-diffusion coupling indeed results in enhanced rotational diffusion at long times for an axisymmetric particle under shear. In a later paper (Leahy et al., 2015), a continuum theory is developed to calculate the time-dependent orientation distribution for non-spherical axisymmetric particles confined to rotate in the velocity-gradient plane, in the limit of weak diffusion or large Péclet number (see sec. 3.1 for the definition). In this limit, a coordinate transformation is discovered and used to map the orientation dynamics to a diffusion equation, which ultimately allowed the calculation of the long-time rotational dispersion coefficient. Furthermore, a remarkably simple analytic expression is obtained for the dispersion coefficient in the large Péclet number limit. In comparison to the classical Taylor dispersion, Leahy et al. (2015) concluded that their theoretical consideration does not fit nicely under the canonical GTD framework.
In this work, we show that the flow-enhanced rotational transport in the velocity-gradient plane can be treated as a GTD in orientation space. To setup the system for such a treatment, the key step is to consider the unbounded angular displacement $\varphi$ rather than the orientation angle $\phi$, which is bounded to an interval of length $2\pi$. With this, one can then break down the unbounded displacement into an infinite sequence of cells, each of which has a length of $2\pi$. In the language of GTD, one then identifies the cell index $j\in\mathbb{Z}$ as the global coordinate and $\phi$ as the local coordinate. The derived GTD theory works for generic linear flows and for arbitrary Péclet numbers. In the large Péclet number limit, we show that our asymptotic result agrees with that obtained by Leahy et al. (2015) for steady simple shear. Our results from the GTD theory is validated against Brownian dynamics simulations.
In section 2, starting from the Smoluchowski equation governing the orientational dynamics of a spheroidal particle, we develop the GTD formulation for generic linear flows. Following Leahy et al. (2015), the particle is constrained to rotate in the velocity-gradient plane. In section 3, we consider the long-time rotational transport in simple shear and extensional flows. The transport coefficients are calculated using perturbation expansions in both small and large Péclet number limits. The results obtained in these asymptotic limits are compared with numerical solutions of the macrotransport equations and results from Brownian dynamics simulations. Lastly, we conclude the paper in section 4.
@@ 2.1 The Smoluchowski equation
We note that (2) can be treated as a marginalization of the full probability density function $P(\bm{x},\bm{q},t)$ that describes the joint distribution of the particle in both position and orientation space, where $\bm{x}$ is the position vector of the particle center. This full probability is governed by
@@ 2.2 Rotational Taylor dispersion theory
It is cumbersome to work with the unit orientation vector $\bm{q}$ in the consideration of the long-time rotational dispersion because $\bm{q}$ is bounded to the unit sphere. From a micromechanical perspective in considering the stochastic trajectory of a particle, one needs to be able to track the unbounded or cumulative angular displacement. The particle orientation vector is constrained to rotate in the velocity-gradient plane. In two dimensions, the orientation vector can be parameterized as $\bm{q}=\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y}$, where $\bm{e}_{x}$ and $\bm{e}_{y}$ are the unit basis vectors of the Cartesian coordinate system $(x,y)$, and $\phi\in[0,2\pi)$ is the orientation angle. The cumulative angular displacement $\varphi$ that is not bounded to the interval $[0,2\pi)$ can be defined via
@@
We remark that alternative methods exist to quantify the rotational dynamics. In particular, one may extract a long-time dispersion coefficient from an orientational correlation function as a function of time in a Brownian dynamics simulation of the orientational Lagenvin equation of motion (Dhont, 1996; Zwanzig, 2001; Leahy et al., 2013). One can also directly keep track of the unbounded angular displacement $\varphi$ in a Brownian dynamics simulation and infer the long-time transport coefficients (Kämmerer et al., 1997; De Michele & Leporini, 2001; Mazza et al., 2006, 2007; Hunter et al., 2011). Because our aim is to derive a generalized Taylor dispersion theory from a continuum (Smoluchowski) perspective, such methods are not pursued here. In Leahy et al. (2015), a coordinate transformation is discovered and used to map the orientational dynamics to a diffusion equation, which ultimately leads to a closed-form asymptotic solution to the long-time rotational diffusivity in the high shear rate limit.
@@
where the angular velocity $\Omega(\phi)$ depends on the orientation angle. It is clear that (5) remains unchanged in terms of the unbounded coordinate $\varphi$. Noting that $\Psi=\Psi(\varphi,t)=\Psi(j,\phi,t)$, one can rewrite $\Psi$ in terms of the sequence $\Psi_{j}(\phi,t)$, $\forall j\in\mathbb{Z}$. In other words, to locate the particle in the unbounded orientation space, one can first identify the cell index, $j$, in which the particle resides, and then use the local angular position $\phi$ within this cell. In the language of generalized Taylor dispersion, $\phi$ is identified as the local coordinate whereas the cell index $j$ is the global coordinate (Frankel & Brenner, 1989; Brenner & Edwards, 1993). In other words, the long-time diffusive dynamics holds only when the particles have traversed many cells.
Because the $\varphi$ space is unbounded, it is more convenient to work in Fourier space. In the following, we make use of the flux-averaging approach of Brady and coworkers (Morris & Brady, 1996; Zia & Brady, 2010; Takatori & Brady, 2014, 2017; Burkholder & Brady, 2017; Peng & Brady, 2020). We note that the original approach was developed for the transport of particles in unbounded domains where the global coordinate is continuous (e.g., the longitudinal coordinate along a flat channel). Recently, Barakat & Takatori (2023) extended this approach to accommodate the transport and dispersion in an oscillating array of harmonic traps, where the global coordinate is the discrete cell index of the infinite lattice. (An equivalent real-space approach may be taken where one make use of the method of moments (Brenner, 1980; Alonso-Matilla et al., 2019). ) In the current problem, we have a one-dimensional lattice of unit cells. Following Barakat & Takatori (2023), we introduce the semidiscrete Fourier transform (Trefethen, 2000)
@@
where $k$ is the wavenumber, and $i$ ($i^{2}=-1$) is the imaginary unit. Notice that the transform is from $j$ to $k$, and the local coordinate $\phi$ is unchanged. In Fourier space, the Smoluchowski equation becomes
@@
where $\hat{\Psi}=\hat{\Psi}(k,\phi,t)$ is the Fourier transform of $\Psi$. The cell-averaged distribution,
@@
which is obtained by averaging (7). In writing (9), we have invoked the periodic boundary condition on the local coordinate $\phi$ (Barakat & Takatori, 2023). One can relate $\hat{\Psi}$ to its average by defining the structure function $\hat{G}$ such that $\hat{\Psi}(k,\phi,t)=\big{\langle}\hat{\Psi}\big{\rangle}(k,t)\hat{G}(k,\phi,t)$. The structure function is normalized
@@
To derive an effective advection-diffusion equation for $\big{\langle}\hat{\Psi}\big{\rangle}$, we first take a small wavenumber expansion of $\hat{G}$, giving
@@
Notice that (12a) is an effective advection-diffusion equation in Fourier space, with the effective rotational drift and rotational dispersion coefficient given in (12b).
@@
The average displacement field vanishes: $\langle b\rangle=0$. Equations (2.2), (14), and (15) are the main results of this paper.
It follows that for a particle undergoing a steady rotation ($\Omega=const.$), $g=1$ and $b=0$, which implies that $\Omega^{\mathrm{eff}}=\Omega$ and $D^{\mathrm{eff}}=D_{R}$. As a result, a nonuniform or $\bm{q}$-dependent angular velocity is required to achieve a long-time dispersivity that is potentially different from the bare diffusivity $D_{R}$. To calculate the average drift, one needs to solve (14) and then take the average of $\Omega g$. With the solution of (14), one can solve (15) and then use (12b) to calculate the dispersion coefficient.
@@ 3.1 Simple shear
Consider the simple shear flow given by $\bm{u}=\dot{\gamma}y\bm{e}_{x}$, where $\bm{e}_{x}$ is the unit basis vector in the $x$ direction of the Cartesian coordinate system $(x,y,z)$, $\dot{\gamma}$ is the shear rate. The problem is non-dimensionalized using the time scale $\tau_{R}=1/D_{R}$. Two dimensionless groups dictate the behavior of the problem. The first is a Péclet number, $Pe=\dot{\gamma}\tau_{R}$, which compares the time scale of the flow to that of rotational diffusion. The second parameter is the Bretherton constant that characterizes the aspect ratio of the spheroid (Bretherton, 1962). The non-dimensional (scaled by $\tau_{R}$) angular velocity is
@@
For spherical particles, $B=0$, and the angular velocity is a constant, $\Omega\tau_{R}=-Pe/2$. In this case, $g=1$ and $\Omega^{\mathrm{eff}}=\Omega$. From (15), we readily obtain $b=0$ and $D^{\mathrm{eff}}=D_{R}$. Because the angular velocity is a constant, the average drift is simply the angular velocity of the flow and the flow does not affect the dispersion coefficient. Similar to the classical Taylor dispersion in which a spatially non-uniform advection is present, an orientation-dependent angular velocity is required to have potentially flow-enhanced dispersion.
@@
where $C=-36175872+51011584Pe^{2}-919040Pe^{4}-47328Pe^{6}-170Pe^{8}+Pe^{10}$, and odd terms vanish.
@@
In figure 1, we plot the the average drift scaled by the drift of a sphere and the dispersion coefficient as a function of $Pe$ for several values of $B$. The scaled drift is shown in figure 1(a) and the non-dimensional dispersion coefficient is plotted in figure 1(b). The truncated series solution is shown in dashed lines. The circles in figure 1 are results obtained by solving (14) and (15) at steady state using a Fourier collocation method. The squares are from Brownian dynamics simulations of the orientational Langevin equation. In 2D, the Langevin equation (dimensional) is written as $\text{d}\varphi/dt=\Omega+\sqrt{2D_{R}}\xi$, where $\xi$ is a white-noise process satisfying $\overline{\xi(t)}=0$ and $\overline{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$. Here, the overhead bar denotes an ensemble average and $\delta$ is the delta function. We remark that in the Langevin equation, the unbounded angular coordinate $\varphi$ is used in order to calculate the mean and mean-squared angular displacements, from which the drift and dispersion coefficient can be obtained. The full numerical solutions (circles) of (14) and (15) agree with the results from BD (squares), which validates our theory.
@@
For spheres, $B=0$, and the drift is equal to the constant angular velocity. As $B$ increases, the drift decreases compared to that of the sphere because the alignment term due to the rate-of-strain becomes more important. In the limit $Pe\to 0$, the drift of nonspherical particles approaches that of spheres, $\Omega^{\mathrm{eff}}/\Omega^{\mathrm{eff}}_{B=0}\to 1$. The reduction of the scaled drift occurs at nonzero $Pe$ and is most prominent for large $Pe$ where the scaled drift asymptotes to a plateau as $Pe\to\infty$. For non-spherical particles, we observe a shear-enhanced angular dispersion as shown in figure 1(b). Similar to the classical Taylor dispersion in linear position space, the enhanced angular dispersion is a result of the coupling between nonuniform advection and diffusion. For spheres, the angular velocity is constant, and no shear-enhanced dispersion is observed. When $B>0$, the dispersion coefficient increases monotonically as a function of $Pe$ until it asymptotes to a plateau at large $Pe$. In dimensional terms, we have $D^{\mathrm{eff}}/D_{R}=O(Pe^{0})$ as $Pe\to\infty$, which is different from the classical Taylor dispersion of Brownian solutes in Poiseuille flow where $D^{\mathrm{eff}}/D_{T}\sim Pe^{2}$ as $Pe\to\infty$. For dispersion in position space, $D_{T}$ is the translational diffusivity and $Pe=UL/D_{T}$ with $U$ the characteristic fluid velocity, $L$ the characteristic width of the channel. We further note that Brownian particles in unbounded shear flows exhibit anomalous diffusion in position space (San Miguel & Sancho, 1979; Foister & Van De Ven, 1980; Katayama & Terauti, 1996; Orihara & Takikawa, 2011; Takikawa & Orihara, 2012).
@@
To understand the behavior of the system at large $Pe$, we consider a perturbation expansion in powers of $1/Pe$,
@@
Inserting the expansion (21a) into (14), one can solve the resulting equations order by order. We derive after some algebra
@@
We note that in obtaining the preceding solutions, the conservation conditions $\langle g_{0}\rangle=1$ and $\langle g_{1}\rangle=0$ are enforced. From (3.1), we obtain
@@
As $Pe\to\infty$, the scaled drift approaches a finite value that depends on $B$. Due to symmetry, the function $g_{1}$ does not contribute to the drift. The correction is $O(1/Pe^{2})$, which comes from $g_{2}$ in the expansion. In figure 2(a), the leading-order asymptotic solution $g_{0}$ is plotted as a function of $\phi$ for $B=0.6$. The full numerical solution of (14) at $Pe=1000$ and $B=0.6$ is also shown in figure 2(a). In figure 4(a), we plot the leading-order drift given in (23) as a function of $B$ and the numerical solution of the macrotransport equations for $Pe=1000$. Good agreement between the asymptotic and numerical solutions is observed.
@@
Substituting (21b) and (3.1) into (15), we obtain at $O(1)$
@@
In figure 2(b), the leading-order asymptotic solution $b_{0}$ is plotted as a function of $\phi$ for $B=0.6$. The full numerical solution of (15) at $Pe=1000$ and $B=0.6$ is also shown in figure 2(b). Good agreement between the asymptotic and numerical solutions is observed. Because of the symmetry of $\Omega$ and $b_{0}$, the average $\langle\Omega b_{0}\rangle$ vanishes.
To obtain the first nonzero term of shear-induced dispersion, we therefore need to consider the $O(1/Pe)$ solution to $b$. At $O(1/Pe)$, the displacement field equation is given by
@@
In figure 4(b), we compare the asymptotic solution given in (32) with the full numerical solution of the macrotransport equations for $Pe=1000.$ We note that (32) agrees with the result obtained by Leahy et al. (2015) where a coordinate transformation is employed to map the orientational dynamics to a diffusion equation in the large $Pe$ limit. In contrast to their approach, the current theory follows closely the GTD framework and allows us to calculate both the average drift and effective dispersion for arbitrary Péclet numbers in generic linear flows.

### v2

@@ Abstract
The coupling between advection and diffusion in position space can often lead to enhanced mass transport compared with diffusion without flow. An important framework used to characterize the long-time diffusive transport in position space is the generalized Taylor dispersion theory. In contrast, the dynamics and transport in orientation space remains less developed. In this work we develop a rotational Taylor dispersion theory that characterizes the long-time orientational transport of a spheroidal particle in linear flows that is constrained to rotate in the velocity-gradient plane. Similar to Taylor dispersion in position space, the orientational distribution of axisymmetric particles in linear flows at long times satisfies an effective advection-diffusion equation in orientation space. Using this framework, we then calculate the long-time average angular velocity and dispersion coefficient for both simple shear and extensional flows. Analytic expressions for the transport coefficients are derived in several asymptotic limits including nearly spherical particles, weak flow and strong flow. Our analysis shows that at long times the effective rotational dispersion is enhanced in simple shear and suppressed in extensional flow. The asymptotic solutions agree with full numerical solutions of the derived macrotransport equations and results from Brownian dynamics simulations. Our results show that the interplay between flow-induced rotations and Brownian diffusion can fundamentally change the long-time transport dynamics.
@@
colloids, dispersion, microscale transport
@@ 1 Introduction
Transport and mixing of solutes or particles in the presence of hydrodynamic flows are important for various biological and industrial processes. For micron-sized particles immersed in flows, the coupling between advection and diffusion can often lead to enhanced mass transport as compared with diffusion without flow. A classical example of such a coupling effect is the Taylor dispersion of Brownian solutes in pressure-driven channel flows (Taylor, 1953, 1954a, 1954b; Aris, 1956). Brownian motion allows the solute particles to migrate across streamlines and then be advected downstream with different velocities. At long times, the coupling between transverse diffusion and longitudinal advection gives rise to diffusive transport of the solutes with an effective longitudinal dispersion coefficient that can be much larger than the bare diffusivity of the solute particle. Since the work of Taylor (1953), a generalized Taylor dispersion (GTD) framework (Frankel & Brenner, 1989) has been developed to accommodate a wide range of transport problems including complex geometries, chemical reactions, spatial and/or time periodicity and active particles (Brenner, 1980; Shapiro & Brenner, 1986, 1987, 1990; Morris & Brady, 1996; Hill & Bees, 2002; Bearon, 2003; Manela & Frankel, 2003; Zia & Brady, 2010; Takatori & Brady, 2014; Burkholder & Brady, 2017; Alonso-Matilla et al., 2019; Jiang & Chen, 2019; Peng & Brady, 2020). More recently, longitudinal dispersion of elongated Brownian rods in a two-dimensional Poiseuille flow has been considered (Kumar et al., 2021; Khair, 2022).
@@
Using experiments and particle-based simulations, Leahy et al. (2013) showed that advection-diffusion coupling indeed results in enhanced rotational diffusion at long times for an axisymmetric particle under shear. In a later paper (Leahy et al., 2015), a continuum theory is developed to calculate the time-dependent orientation distribution for non-spherical axisymmetric particles confined to rotate in the velocity-gradient plane, in the limit of weak diffusion or large Péclet number (see sec. 3.1 for the definition). In this limit, a coordinate transformation is discovered and used to map the orientation dynamics to a diffusion equation, which ultimately allowed the calculation of the long-time rotational dispersion coefficient. Furthermore, a remarkably simple analytic expression is obtained for the dispersion coefficient in the large-Péclet-number limit. In comparison to the classical Taylor dispersion, Leahy et al. (2015) concluded that their theoretical consideration does not fit nicely under the canonical GTD framework.
In this work we show that the flow-enhanced rotational transport in the velocity-gradient plane can be treated as a GTD in orientation space. To setup the system for such a treatment, the key step is to consider the unbounded angular displacement $\varphi$ rather than the orientation angle $\phi$, which is bounded to an interval of length $2\pi$. With this, one can then break down the unbounded displacement into an infinite sequence of cells, each of which has a length of $2\pi$. In the language of GTD, one then identifies the cell index $j\in\mathbb{Z}$ as the global coordinate and $\phi$ as the local coordinate. The derived GTD theory works for generic linear flows and for arbitrary Péclet numbers. In the large-Péclet-number limit, we show that our asymptotic result agrees with that obtained by Leahy et al. (2015) for steady simple shear. Our results from the GTD theory is validated against Brownian dynamics (BD) simulations.
In §$\S$ 2, starting from the Smoluchowski equation governing the orientational dynamics of a spheroidal particle, we develop the GTD formulation for generic linear flows. Following Leahy et al. (2015), the particle is constrained to rotate in the velocity-gradient plane. In §$\S$ 3, we consider the long-time rotational transport in simple shear and extensional flows. The transport coefficients are calculated using perturbation expansions in both small- and large-Péclet-number limits. The results obtained in these asymptotic limits are compared with numerical solutions of the macrotransport equations and results from BD simulations. Lastly, we conclude the paper in §$\S$ 4.
@@ 2.1 The Smoluchowski equation
We note that (2) can be treated as a marginalization of the full probability density function $P(\bm{x},\bm{q},t)$ that describes the joint distribution of the particle in both position and orientation space, where $\bm{x}$ is the position vector of the particle centre. This full probability is governed by
@@ 2.2 Rotational Taylor dispersion theory
It is cumbersome to work with the unit orientation vector $\bm{q}$ in the consideration of the long-time rotational dispersion because $\bm{q}$ is bounded to the unit sphere (Kämmerer et al., 1997; De Michele & Leporini, 2001; Mazza et al., 2006, 2007; Hunter et al., 2011). From a micromechanical perspective in considering the stochastic trajectory of a particle, one needs to be able to track the unbounded or cumulative angular displacement. The particle orientation vector is constrained to rotate in the velocity-gradient plane (Leahy et al., 2013, 2015). In two dimensions, the orientation vector can be parameterized as $\bm{q}=\cos\phi\bm{e}_{x}+\sin\phi\bm{e}_{y}$, where $\bm{e}_{x}$ and $\bm{e}_{y}$ are the unit basis vectors of the Cartesian coordinate system $(x,y)$, and $\phi\in[0,2\pi)$ is the orientation angle. The cumulative angular displacement $\varphi$ that is not bounded to the interval $[0,2\pi)$ can be defined via
@@
We remark that alternative methods exist to quantify the rotational dynamics. In particular, one may extract a long-time dispersion coefficient from an orientational correlation function as a function of time in a BD simulation of the orientational Langevin equation of motion (Dhont, 1996; Zwanzig, 2001; Leahy et al., 2013). One can also directly keep track of the unbounded angular displacement $\varphi$ in a BD simulation and infer the long-time transport coefficients (Kämmerer et al., 1997; De Michele & Leporini, 2001; Mazza et al., 2006, 2007; Hunter et al., 2011). Because our aim is to derive a GTD theory from a continuum (Smoluchowski) perspective, such methods are not pursued here. In Leahy et al. (2015) a coordinate transformation is discovered and used to map the orientational dynamics to a diffusion equation, which ultimately leads to a closed-form asymptotic solution to the long-time rotational diffusivity in the high shear rate limit.
@@
where the angular velocity $\Omega(\phi)$ depends on the orientation angle. It is clear that (5) remains unchanged in terms of the unbounded coordinate $\varphi$. Noting that $\Psi=\Psi(\varphi,t)=\Psi(j,\phi,t)$, one can rewrite $\Psi$ in terms of the sequence $\Psi_{j}(\phi,t)$, $\forall j\in\mathbb{Z}$. In other words, to locate the particle in the unbounded orientation space, one can first identify the cell index, $j$, in which the particle resides, and then use the local angular position $\phi$ within this cell. In the language of GTD, $\phi$ is identified as the local coordinate whereas the cell index $j$ is the global coordinate (Frankel & Brenner, 1989; Brenner & Edwards, 1993). In other words, the long-time diffusive dynamics holds only when the particles have traversed many cells.
Because the $\varphi$ space is unbounded, it is more convenient to work in Fourier space. In the following, we make use of the flux-averaging approach of Brady and coworkers (Morris & Brady, 1996; Zia & Brady, 2010; Takatori & Brady, 2014, 2017; Burkholder & Brady, 2017; Peng & Brady, 2020). We note that the original approach was developed for the transport of particles in unbounded domains where the global coordinate is continuous (e.g., the longitudinal coordinate along a flat channel). Recently, Barakat & Takatori (2023) extended this approach to accommodate the transport and dispersion in an oscillating array of harmonic traps, where the global coordinate is the discrete cell index of the infinite lattice. (An equivalent real-space approach may be taken where one makes use of the method of moments; see, e.g. Brenner 1980; Alonso-Matilla et al. 2019. ) In the current problem, we have a one-dimensional lattice of unit cells. Following Barakat & Takatori (2023), we introduce the semi-discrete Fourier transform (Trefethen, 2000)
@@
where $k$ is the wavenumber and $i$ ($i^{2}=-1$) is the imaginary unit. Note that the transform is from $j$ to $k$, and the local coordinate $\phi$ is unchanged. In Fourier space, the Smoluchowski equation becomes
@@
where $\hat{\Psi}=\hat{\Psi}(k,\phi,t)$ is the Fourier transform of $\Psi$. We note that in (7) the local gradient is written in real space whereas the global gradient appears in powers of $k$. The cell-averaged distribution,
@@
which is obtained by averaging (7). In writing (9) we have invoked the periodic boundary condition on the local coordinate $\phi$ (Barakat & Takatori, 2023). One can relate $\hat{\Psi}$ to its average by defining the structure function $\hat{G}$ such that $\hat{\Psi}(k,\phi,t)=\big{\langle}\hat{\Psi}\big{\rangle}(k,t)\hat{G}(k,\phi,t)$. The structure function is normalized, i.e.
@@
To derive an effective advection-diffusion equation for $\big{\langle}\hat{\Psi}\big{\rangle}$, we first take a small wavenumber expansion of $\hat{G}$ (Morris & Brady, 1996; Zia & Brady, 2010), giving
@@
Note that (12a) is an effective advection-diffusion equation in Fourier space, with the effective rotational drift and rotational dispersion coefficient given in (12b).
@@
The average displacement field vanishes: $\langle b\rangle=0$. Equations (12), (14) and (15) are the main results of this paper.
It follows that for a particle undergoing a steady rotation ($\Omega=const.$), $g=1$ and $b=0$, which implies that $\Omega^{\mathrm{eff}}=\Omega$ and $D^{\mathrm{eff}}=D_{R}$. As a result, a non-uniform or $\bm{q}$-dependent angular velocity is required to achieve a long-time dispersivity that is potentially different from the bare diffusivity $D_{R}$. To calculate the average drift, one needs to solve (14) and then take the average of $\Omega g$. With the solution of (14), one can solve (15) and then use (12b) to calculate the dispersion coefficient.
@@ 3.1 Simple shear
Consider the simple shear flow given by $\bm{u}=\dot{\gamma}y\bm{e}_{x}$, where $\bm{e}_{x}$ is the unit basis vector in the $x$ direction of the Cartesian coordinate system $(x,y,z)$, $\dot{\gamma}$ is the shear rate. The problem is non-dimensionalized using the time scale $\tau_{R}=1/D_{R}$. Two dimensionless groups dictate the behaviour of the problem. The first is a Péclet number, $Pe=\dot{\gamma}\tau_{R}$, which compares the time scale of the flow to that of rotational diffusion. The second parameter is the Bretherton constant that characterizes the aspect ratio of the spheroid (Bretherton, 1962). The non-dimensional (scaled by $\tau_{R}$) angular velocity is
@@
For spherical particles, $B=0$, and the angular velocity is a constant, $\Omega\tau_{R}=-Pe/2$. In this case, $g=1$ and $\Omega^{\mathrm{eff}}=\Omega$. From (15), we readily obtain $b=0$ and $D^{\mathrm{eff}}=D_{R}$. Because the angular velocity is a constant, the average drift is simply the angular velocity of the flow and the flow does not affect the dispersion coefficient. Similar to the classical Taylor dispersion in which a spatially non-uniform advection is present, an orientation-dependent angular velocity is required to have potentially flow-enhanced dispersion (Leahy et al., 2013, 2015).
@@
where $C=-36175872+51011584Pe^{2}-919040Pe^{4}-47328Pe^{6}-170Pe^{8}+Pe^{10}$ and odd terms vanish.
@@
In figure 1 we plot the average drift scaled by the drift of a sphere and the dispersion coefficient as a function of $Pe$ for several values of $B$. The scaled drift is shown in figure 1(a) and the non-dimensional dispersion coefficient is plotted in figure 1(b). The truncated series solution is shown by dashed lines. The circles in figure 1 are results obtained by solving (14) and (15) at steady state using a Fourier collocation method. The squares are from BD simulations of the orientational Langevin equation. In two dimensions, the Langevin equation (dimensional) is written as $\text{d}\varphi/dt=\Omega+\sqrt{2D_{R}}\xi$, where $\xi$ is a white-noise process satisfying $\overline{\xi(t)}=0$ and $\overline{\xi(t)\xi(t^{\prime})}=\delta(t-t^{\prime})$. Here, the overhead bar denotes an ensemble average and $\delta$ is the delta function. We remark that in the Langevin equation, the unbounded angular coordinate $\varphi$ is used in order to calculate the mean and mean-squared angular displacements, from which the drift and dispersion coefficient can be obtained. The full numerical solutions (circles) of (14) and (15) agree with the results from BD (squares), which validates our theory.
@@
For spheres, $B=0$, and the drift is equal to the constant angular velocity. As $B$ increases, the drift decreases compared with that of the sphere because the alignment term due to the rate of strain becomes more important. In the limit $Pe\to 0$, the drift of non-spherical particles approaches that of spheres, $\Omega^{\mathrm{eff}}/\Omega^{\mathrm{eff}}_{B=0}\to 1$. The reduction of the scaled drift occurs at non-zero $Pe$ and is most prominent for large $Pe$ where the scaled drift asymptotes to a plateau as $Pe\to\infty$. For non-spherical particles, we observe a shear-enhanced angular dispersion as shown in figure 1(b). Similar to the classical Taylor dispersion in linear position space, the enhanced angular dispersion is a result of the coupling between non-uniform advection and diffusion. For spheres, the angular velocity is constant and no shear-enhanced dispersion is observed. When $B>0$, the dispersion coefficient increases monotonically as a function of $Pe$ until it asymptotes to a plateau at large $Pe$. In dimensional terms, we have $D^{\mathrm{eff}}/D_{R}=O(Pe^{0})$ as $Pe\to\infty$, which is different from the classical Taylor dispersion of Brownian solutes in Poiseuille flow where $D^{\mathrm{eff}}/D_{T}\sim Pe^{2}$ as $Pe\to\infty$. For dispersion in position space, $D_{T}$ is the translational diffusivity and $Pe=UL/D_{T}$ with $U$ the characteristic fluid velocity and $L$ the characteristic width of the channel. We further note that Brownian particles in unbounded shear flows exhibit anomalous diffusion in position space (San Miguel & Sancho, 1979; Foister & Van De Ven, 1980; Krishnan & Leighton, 1992; Katayama & Terauti, 1996; Orihara & Takikawa, 2011; Takikawa & Orihara, 2012).
@@
To understand the behaviour of the system at large $Pe$, we consider a perturbation expansion in powers of $1/Pe$,
@@
Inserting the expansion (21a) into (14), one can solve the resulting equations order by order. We derive after some algebra that
@@
We note that in obtaining the preceding solutions, the conservation conditions $\langle g_{0}\rangle=1$ and $\langle g_{1}\rangle=0$ are enforced. From (22), we obtain
@@
As $Pe\to\infty$, the scaled drift approaches a finite value that depends on $B$. Due to symmetry, the function $g_{1}$ does not contribute to the drift. The correction is $O(1/Pe^{2})$, which comes from $g_{2}$ in the expansion. In figure 2(a) the leading-order asymptotic solution $g_{0}$ is plotted as a function of $\phi$ for $B=0.6$. The full numerical solution of (14) at $Pe=1000$ and $B=0.6$ is also shown in figure 2(a). In figure 4(a) we plot the leading-order drift given in (23) as a function of $B$ and the numerical solution of the macrotransport equations for $Pe=1000$. Good agreement between the asymptotic and numerical solutions is observed.
@@
Substituting (21b) and (22) into (15), we obtain at $O(1)$
@@
In figure 2(b) the leading-order asymptotic solution $b_{0}$ is plotted as a function of $\phi$ for $B=0.6$. The full numerical solution of (15) at $Pe=1000$ and $B=0.6$ is also shown in figure 2(b). Good agreement between the asymptotic and numerical solutions is observed. Because of the symmetry of $\Omega$ and $b_{0}$, the average $\langle\Omega b_{0}\rangle$ vanishes.
To obtain the first non-zero term of shear-induced dispersion, we therefore need to consider the $O(1/Pe)$ solution to $b$. At $O(1/Pe)$, the displacement field equation is given by
@@
In figure 4(b) we compare the asymptotic solution given in (32) with the full numerical solution of the macrotransport equations for $Pe=1000$. We note that (32) agrees with the result obtained by Leahy et al. (2015) where a coordinate transformation is employed to map the orientational dynamics to a diffusion equation in the large-$Pe$ limit. In contrast to their approach, the current theory follows closely the GTD framework and allows us to calculate both the average drift and effective dispersion for arbitrary Péclet numbers in generic linear flows.